# Low-rank orthogonalization for large-scale matrix optimization with applications to foundation model training

### v1

@@ 3.2 Low-rank matrix-signed gradient descent
In this subsection, we propose low-rank matrix-signed gradient descent methods, including a fixed-rank variant and a safeguarded variant with adaptive ranks. Both variants use low-rank orthogonalization as a subroutine.
Algorithm 2 Low-rank matrix-signed gradient descent
@@
  Call Algorithm 1 (see Section 3.1) with $(M,r)=(\nabla f(X^{k}),r)$ to obtain an approximate matrix sign
  $M^{k}_{O}$, such that $M^{k}_{O}=\mathrm{msgn}(M^{k}_{Q})$, where $M^{k}_{Q}$ is a low-rank approximation for $\nabla f(X^{k})$.
  Update the next iterate:
$X^{k+1}=X^{k}-\eta_{k}M^{k}_{O}.$
@@
At each iteration $k\geq 0$, the fixed-rank variant invokes Algorithm 1 to compute $M^{k}_{O}=\mathrm{msgn}(M_{Q}^{k})$, where $M_{Q}^{k}$ is a low-rank approximation of $\nabla f(X^{k})$. Then, the next iterate $X^{k+1}$ is obtained by performing a line search update from $X^{k}$ along the matrix-signed direction $M^{k}_{O}$ with a suitable step size. Details of this method are presented in Algorithm 2.
@@
Suppose that Assumption 1 holds. Let $\{(X^{k},M_{O}^{k})\}$ be the sequence generated by Algorithm 2 with $M_{O}^{k}=\mathrm{msgn}(M_{Q}^{k})$ for all $k\geq 0$ and step sizes $\{\eta_{k}\}$ given by
$\eta_{k}=\frac{1}{(k+1)^{1/2}}\qquad\forall k\geq 0.$ (7)
Then, it holds that for all $K\geq 3$,
@@
where $L_{*}$ are defined in Assumption 1.
@@
From Theorem 2, we observe that when $\{(X^{k},M_{Q}^{k})\}$ satisfies $\sum_{k=0}^{K-1}\|\nabla f(X^{k})-M_{Q}^{k}\|_{*}(k+1)^{-1/2}=\widetilde{\mathcal{O}}(1)$, it holds that $\min_{0\leq k\leq K-1}\{\|\nabla f(X^{k})\|_{*}\}=\widetilde{\mathcal{O}}(K^{-1/2})$, which matches, up to a logarithmic factor, the well-established optimal convergence rate for nonconvex optimization (e.g., see [13]). Moreover, if the gradients $\{\nabla f(X^{k})\}$ have low rank when $k$ sufficiently large (as is often the case during deep neural network training [57]), we can tune the rank parameter $r$ such that it is close to the effective rank of the gradients, making the sequence $\{\|\nabla f(X^{k})-M_{Q}^{k}\|_{*}\}$ to remain close to zero.
We next describe a safeguarded low-rank matrix-signed gradient descent method with adaptively updated ranks. At each iteration $k\geq 0$, this method invokes Algorithm 1 with $(M,r)=(\nabla f(X^{k}),r_{k})$ to obtain $M^{k}_{O}=\mathrm{msgn}(M_{Q}^{k})$, where $M_{Q}^{k}$ is a low-rank approximation of $\nabla f(X^{k})$ such that the approximation error satisfies (9). Then, the next iterate $X^{k+1}$ is obtained by performing a line search update from $X^{k}$ along the matrix-signed direction $M^{k}_{O}$, with a suitable step size. Details of this method are provided in Algorithm 3.
It is noteworthy that the approximation error in (9) holds for some $r_{k}\leq\varrho$ because when $r_{k}=\varrho$, the error $\|M^{k}_{Q}-\nabla f(X^{k})\|$ is zero. When the matrix is low-rank or has a small effective rank, $r_{k}$ can be chosen to be much smaller than $\varrho$. In practice, one can gradually increase the trial ranks to find $r_{k}$ such that (9) holds.
Algorithm 3 Safeguarded low-rank matrix-signed gradient descent
@@
  Call Algorithm 1 (see Section 3.1) with $(M,r)=(\nabla f(X^{k}),r_{k})$ to obtain an approximate matrix sign
  $M^{k}_{O}$ such that $M_{O}^{k}=\mathrm{msgn}(M_{Q}^{k})$ and
@@
  Update the next iterate:
$X^{k+1}=X^{k}-\eta_{k}M^{k}_{O}.$
@@
Let $\{X^{k}\}$ be generated by Algorithm 3 with input parameters $\{(\eta_{k},\delta_{k})\}$ given by
$\eta_{k}=\delta_{k}=\frac{1}{(k+1)^{1/2}}\qquad\forall k\geq 0.$ (11)
Then, for any $\epsilon\in(0,1)$, it holds that $\min_{0\leq k\leq K-1}\{\|\nabla f(X^{k})\|_{*}\}\leq\epsilon$ for all $K$ satisfying
@@
From Theorem 3, we observe that Algorithm 3 achieves an iteration complexity of $\widetilde{\mathcal{O}}(\epsilon^{-2})$ for finding an $\epsilon$-NSP of (1). This complexity bound matches, up to a polylogarithmic factor, the lower complexity bound as established in [13].
@@ 3.3 Low-rank Muon
In this subsection, we propose a low-rank variant of Muon [28], and analyze its iteration complexity under heavy-tailed noise.
This method follows a framework similar to Muon [28], but instead of directly computing the matrix sign of the momentum update, it computes only its low-rank approximation. Specifically, this method generates three sequences $\{X^{k}\}$, $\{M^{k}\}$, and $\{M^{k}_{O}\}$. At each iteration $k\geq 0$, it first performs a momentum update to generate $M^{k}$ by aggregating stochastic gradients of $f$ evaluated at $X^{0},\ldots,X^{k}$. Then, Algorithm 1 is invoked to obtain $M^{k}_{O}=\mathrm{msgn}(M_{Q}^{k})$, where $M_{Q}^{k}$ is a low-rank approximation of $\nabla f(X^{k})$ such that the approximation error satisfies (13). The next iterate $X^{k+1}$ is obtained by performing a line search update from $X^{k}$ along the matrix-signed direction $M_{O}^{k}$ with a suitable step size. Details of this method are given in Algorithm 4.
@@
  Call Algorithm 1 (see Section 3.1) with $(M,r)=(M^{k},r_{k})$ to obtain an approximate matrix sign
  $M^{k}_{O}$ such that $M_{O}^{k}=\mathrm{msgn}(M^{k}_{Q})$ and
@@
Suppose that Assumptions 1 and 2 hold. Let $f_{\mathrm{low}}$ and $L_{*}$ be given in Assumption 1, and $\sigma$ and $\alpha$ be given in Assumption 2, and define
@@
Let $\{X^{k}\}$ be generated by Algorithm 4 with input parameters $\{(\eta_{k},\theta_{k},\delta_{k})\}$ given by
$\eta_{k}=\frac{1}{(k+1)^{(2\alpha-1)/(3\alpha-2)}},\quad\theta_{k}=\frac{1}{(k+1)^{\alpha/(3\alpha-2)}},\quad\delta_{k}=\frac{1}{(k+1)^{(\alpha-1)/(3\alpha-2)}}\qquad\forall k\geq 0.$ (16)
Then, for any $\epsilon\in(0,1)$, it holds that $\mathbb{E}[\|\nabla f(X^{\iota_{K}})\|_{*}]\leq\epsilon$ for all $K$ satisfying
$K\geq\max\Big{\{}\Big{(}\frac{2(3\alpha-2)U_{\mathrm{mn}}}{(\alpha-1)\epsilon}\ln\Big{(}\frac{2(3\alpha-2)U_{\mathrm{mn}}}{(\alpha-1)\epsilon}\Big{)}\Big{)}^{(3\alpha-2)/(\alpha-1)},3\Big{\}},$
@@
From Theorem 4, one can observe that Algorithm 4 achieves an iteration complexity of $\widetilde{\mathcal{O}}(\epsilon^{-(3\alpha-2)/(\alpha-1)})$ for finding an $\epsilon$-SNSP of problem (1), which matches the optimal dependence on $\epsilon$ in the complexity results for vector-variate stochastic first-order methods under heavy-tailed noise (see, e.g., [24, 56]). To the best of our knowledge, our result is the first to show that the Muon-type algorithm can achieve an iteration complexity with optimal dependence on $\epsilon$ when applied to matrix-variate optimization under heavy-tailed noise.
@@ 4 Numerical experiments
In this section, we present numerical experiments to evaluate the performance of low-rank Muon (Algorithm 4) and compare our method with vanilla Muon, AdamW, and SGD. The experiments are conducted on a synthetic matrix regression problem (Section 4.1) and foundation model pretraining with GPT-2 (Section 4.2) and LLaMA (Section 4.3). All experiments are executed on a server with two NVIDIA A100 GPUs (80 GB).
4.1 Matrix regression
In this subsection, we consider the matrix regression problem:
$\min_{X\in\mathbb{R}^{n\times n}}\Big{\{}f(X)=\frac{1}{2}\|AXB-C\|_{F}^{2}\Big{\}},$ (17)
where $A\in\mathbb{R}^{p\times n}$, $B\in\mathbb{R}^{n\times p}$, and $C\in\mathbb{R}^{p\times p}$ are given data matrices. We simulate low-rank structure for the gradients $\nabla f(X)$ by choosing $n\gg p=100$. For each $n\in\{10^{3},5\times 10^{3},10^{4}\}$, we set $A=U_{A}\Sigma_{A}V_{A}^{T}$ and $B=U_{B}\Sigma_{B}V_{B}^{T}$, where $\Sigma_{A}=\Sigma_{B}=\mathrm{Diag}([2,\ldots,2^{p}])$, and $U_{A}$, $V_{A}$, $U_{B}$, and $V_{B}$ are randomly generated column-orthogonal matrices. We also generate a ground truth solution $X^{*}$ with all entries independently drawn from the standard Gaussian distribution, and set $C=AXB+10^{-3}E$, where all entries of $E$ are independently drawn from the standard Gaussian distribution.
Figure 4: Objective values (top row) and gradient Frobenius norms (bottom row) for all methods during the first $2\times 10^{4}$ iterations. Results for $n=10^{3}$, $5\times 10^{3}$, and $10^{4}$ are shown in the left, middle, and right columns, respectively. The computation times on GPUs are displayed in the legend for each plot.
We apply low-rank Muon, Muon, AdamW, and SGD to solve (17). We implement two low-rank Muon variants with rank parameters 50 and 100, abbreviated as LR-Muon50 and LR-Muon100, respectively. All methods are implemented with full-batch (deterministic) gradients and initialized from the zero matrix. We compare their performance based on the objective values and the Frobenius norm of the gradient for problem (17), evaluated over the first $2\times 10^{4}$ iterations. The algorithmic parameters are selected to suit each method well in terms of computational performance.
For each $n$, we plot the objective values and the gradient Frobenius norms in Figure 4 to illustrate the convergence behavior of all methods. Figure 4 demonstrates that our low-rank Muon achieves performance comparable to vanilla Muon in iteration count, while significantly outperforming both SGD and AdamW. In addition, SGD fails to converge effectively for problem (17). Furthermore, our low-rank Muon demonstrates significantly faster computation times on GPUs compared to other variants, highlighting the advantages of our low-rank orthogonalization procedures for computational efficiency.
Figure 5: Singular value distributions of the Q, K, and V matrices: across layers at iteration 300 (top row) and over training iterations at layer 16 (bottom row) during GPT-2 pretraining.
@@ 4.2 GPT-2 pretraining
In this subsection, we consider pretraining GPT-2 [42], a transformer-based language model. We experiment with GPT-2 models of sizes 60M, 135M, 350M, and 1B on the same three datasets tested in the Muon GitHub repository [28]: FineWeb10B, FineWeb100B, and FineWebEdu10B.
@@
We apply low-rank Muon, Muon, AdamW, and SGD with momentum (SGDM) for GPT-2 pretraining. We implement two low-rank Muon variants with rank parameters 100 and 200, abbreviated as LR-Muon100 and LR-Muon200, respectively. Following the experiments in [27, 31], we use AdamW to train the embedding and head layers for Muon and low-rank Muon. We initialize all methods with the same weights from pretrained GPT-2 models, and terminate all methods after one training epoch, consisting of 5000 iterations. We compare all methods using validation perplexity, a standard metric in foundation model training (e.g., [42, 50, 52]), which is defined as the exponential of the validation loss and serves to amplify performance differences. The algorithmic parameters are carefully tuned to suit each method well in terms of computational performance. More details about the experimental setups including algorithm and GPT-2 model parameters are deferred to Appendix B.
@@
We visualize the top 100 singular values of the momentum updates $\{M^{k}\}$ associated with the Q, K, and V matrices for a 350M GPT-2 model trained by Muon on the FineWeb100B dataset in Figure 5. The first and second rows of the figure display the top 100 singular values across different training iterations and neural network layers, respectively. This visualization partly supports the low-rank nature of the momentum updates.
We present a comparison of the validation perplexity and computational time in Table 1. From this table, we observe that for GPT-2 with a model size of 60M, our low-rank Muon achieves validation perplexity worse than Muon but better than AdamW and SGDM. For GPT-2 models with larger sizes, our low-rank Muon achieves better validation perplexity than all other competing methods. We also observe that the computational time of our low-rank Muon improves upon that of vanilla Muon, though all Muon-type methods remain slower than SGDM and AdamW. These observations show that our low-rank orthogonalization enables obtaining more generalizable GPT-2 models when the model size is large, but may be less effective for smaller models. In addition, our low-rank orthogonalization saves computational time compared to orthogonalization by Newton-Schulz iterations, but the gain is not as significant as those discussed in Sections 3.1 and 4.1, since the forward and backward passes of neural networks account for the majority of the computational time.
We plot the convergence behavior of validation perplexity versus training iterations in Figure 6. From this figure, we observe that for each model size, the competing methods exhibit similar performance patterns across the three datasets. For the model size of 60M, we observe that all Muon-type methods consistently outperforms AdamW and SGDM, while our low-rank Muon performs slightly worse than the vanilla Muon. For the 135M model size, we observe that our low-rank Muon outperforms the vanilla Muon and other training methods. The vanilla Muon performs similarly to AdamW, and all methods show a large improvement compared to SGDM. For the model size of 350M, we see that our low-rank Muon has slightly better performance than AdamW, while outperforming Muon and SGDM by a larger margin. These observations support that our low-rank Muon improves upon Muon in GPT-2 training and is more effective for models with larger sizes.
With all the above observations and discussions, we conclude that our low-rank Muon is a viable alternative to existing training methods and is particularly more effective than other methods for training GPT-2 models with larger sizes.

### v2

@@ 3.2 Low-Rank Matrix-Signed Gradient Descent
In this subsection, we propose low-rank MSGD methods, including a fixed-rank variant and a safeguarded variant with adaptive ranks. Both variants use low-rank orthogonalization as a subroutine.
Algorithm 2 Low-Rank MSGD
@@
  Call Algorithm 1 (see Section 3.1) with $(M,r)=(\nabla f(X^{k}),r)$ to obtain an approximate matrix sign $M^{k}_{O}$, such that $M^{k}_{O}=\mathrm{msgn}(M^{k}_{Q})$, where $M^{k}_{Q}$ is a low-rank approximation for $\nabla f(X^{k})$.
  Update the next iterate: $X^{k+1}=X^{k}-\eta_{k}M^{k}_{O}$.
@@
At each iteration $k\geq 0$, the fixed-rank variant invokes Algorithm 1 to compute $M^{k}_{O}=\mathrm{msgn}(M_{Q}^{k})$, where $M_{Q}^{k}$ is a low-rank approximation of $\nabla f(X^{k})$. Then, the next iterate $X^{k+1}$ is obtained by performing a line search update from $X^{k}$ along the matrix-signed direction $M^{k}_{O}$. Details of this method are presented in Algorithm 2.
@@
Suppose that Assumption 1 holds. Let $\{(X^{k},M_{O}^{k})\}$ be the sequence generated by Algorithm 2 with $M_{O}^{k}=\mathrm{msgn}(M_{Q}^{k})$ for all $k\geq 0$ and step sizes $\{\eta_{k}\}$ given by $\eta_{k}=(k+1)^{-1/2}$ for all $k\geq 0$. Then, it holds that for all $K\geq 3$,
@@
where $L_{*}$ is defined in Assumption 1.
@@
From Theorem 2, we observe that when $\{(X^{k},M_{Q}^{k})\}$ satisfies $\sum_{k=0}^{K-1}\|\nabla f(X^{k})-M_{Q}^{k}\|_{*}(k+1)^{-1/2}=\widetilde{\mathcal{O}}(1)$, it holds that $\min_{0\leq k\leq K-1}\{\|\nabla f(X^{k})\|_{*}\}=\widetilde{\mathcal{O}}(K^{-1/2})$, which matches, up to a logarithmic factor, the well-established optimal convergence rate for nonconvex optimization (e.g., see (Carmon et al., 2020)). Moreover, if the gradients $\{\nabla f(X^{k})\}$ have low rank when $k$ is sufficiently large (as is often the case during deep neural network training (Zhao et al., 2024)), we can tune the rank parameter $r$ to be close to the effective rank of the gradients, causing the sequence $\{\|\nabla f(X^{k})-M_{Q}^{k}\|_{*}\}$ to remain close to zero.
We next describe a safeguarded low-rank MSGD method with adaptively updated ranks. At each iteration $k\geq 0$, this method invokes Algorithm 1 with $(M,r)=(\nabla f(X^{k}),r_{k})$ to obtain $M^{k}_{O}=\mathrm{msgn}(M_{Q}^{k})$, where $M_{Q}^{k}$ is a low-rank approximation of $\nabla f(X^{k})$ such that the approximation error satisfies (7). Then, the next iterate $X^{k+1}$ is obtained by performing a line search update from $X^{k}$ along the matrix-signed direction $M^{k}_{O}$, with a suitable step size. Details of this method are provided in Algorithm 3.
It is noteworthy that the approximation error in (7) holds for some $r_{k}\leq\varrho$ because when $r_{k}=\varrho$, the error $\|M^{k}_{Q}-\nabla f(X^{k})\|_{*}$ is zero. When the matrix is low-rank or has a small effective rank, $r_{k}$ can be chosen to be much smaller than $\varrho$. In practice, one can gradually increase the trial ranks to find $r_{k}$ such that (7) holds.
Algorithm 3 Safeguarded Low-Rank MSGD
@@
  Call Algorithm 1 (see Section 3.1) with $(M,r)=(\nabla f(X^{k}),r_{k})$ to obtain an approximate matrix sign $M^{k}_{O}$ such that $M_{O}^{k}=\mathrm{msgn}(M_{Q}^{k})$ and
@@
  Update the next iterate: $X^{k+1}=X^{k}-\eta_{k}M^{k}_{O}$.
@@
Let $\{X^{k}\}$ be generated by Algorithm 3 with inputs $\{(\eta_{k},\delta_{k})\}$ given by $\eta_{k}=\delta_{k}={(k+1)^{-1/2}}$ for all $k\geq 0$. Then, for any $\epsilon\in(0,1)$, it holds that $\min_{0\leq k\leq K-1}\{\|\nabla f(X^{k})\|_{*}\}\leq\epsilon$ for all $K$ satisfying
@@
From Theorem 3, we observe that Algorithm 3 achieves an iteration complexity of $\widetilde{\mathcal{O}}(\epsilon^{-2})$ for finding an $\epsilon$-NSP of (1). This complexity bound matches, up to a polylogarithmic factor, the lower complexity bound as established in (Carmon et al., 2020).
@@ 3.3 Low-Rank Muon
In this subsection, we propose a low-rank variant of Muon (Jordan et al., ), and analyze its iteration complexity under heavy-tailed noise.
This method follows a framework similar to Muon (Jordan et al., ), but instead of directly computing the matrix sign of the momentum update, it computes only its low-rank approximation. Specifically, at each iteration of this method, it first performs a momentum update to generate $M^{k}$ by aggregating stochastic gradients of $f$ evaluated at $X^{0},\ldots,X^{k}$. Then, Algorithm 1 is invoked to obtain $M^{k}_{O}=\mathrm{msgn}(M_{Q}^{k})$, where $M_{Q}^{k}$ is a low-rank approximation of $\nabla f(X^{k})$ such that the approximation error satisfies (10). The next iterate $X^{k+1}$ is obtained by performing a line search update from $X^{k}$ along the matrix-signed direction $M_{O}^{k}$ with a suitable step size. Details of this method are given in Algorithm 4.
@@
  Call Algorithm 1 (see Section 3.1) with $(M,r)=(M^{k},r_{k})$ to obtain an approximate matrix sign $M^{k}_{O}$ such that $M_{O}^{k}=\mathrm{msgn}(M^{k}_{Q})$ and
@@
Suppose that Assumptions 1 and 2 hold. Let $f_{\mathrm{low}}$ and $L_{*}$ be given in Assumption 1, and $\sigma$ and $\alpha$ be given in Assumption 2. Define
@@
Let $\{X^{k}\}$ be generated by Algorithm 4 with input parameters $\{(\eta_{k},\theta_{k},\delta_{k})\}$ given by $\eta_{k}=(k+1)^{-(2\alpha-1)/(3\alpha-2)}$, $\theta_{k}=(k+1)^{-\alpha/(3\alpha-2)}$, and $\delta_{k}=(k+1)^{-(\alpha-1)/(3\alpha-2)}$ for all $k\geq 0$. Then, for any $\epsilon\in(0,1)$, it holds that $\mathbb{E}[\|\nabla f(X^{\iota_{K}})\|_{*}]\leq\epsilon$ for all $K$ satisfying
$K\geq\max\Big\{\Big(\frac{2(3\alpha-2)U_{\mathrm{mn}}}{(\alpha-1)\epsilon}\ln\Big(\frac{2(3\alpha-2)U_{\mathrm{mn}}}{(\alpha-1)\epsilon}\Big)\Big)^{\frac{3\alpha-2}{\alpha-1}},3\Big\},$
@@
From Theorem 4, one can observe that Algorithm 4 achieves an iteration complexity of $\widetilde{\mathcal{O}}(\epsilon^{-(3\alpha-2)/(\alpha-1)})$ for finding an $\epsilon$-SNSP of problem (1), which matches the optimal dependence on $\epsilon$ in the complexity results for vector-variate stochastic first-order methods under heavy-tailed noise (see, e.g., (He et al., 2025; Zhang et al., 2020)). To the best of our knowledge, our result is the first to show that the inexact Muon-type algorithms can achieve an iteration complexity with optimal dependence on $\epsilon$ when applied to matrix-variate optimization under heavy-tailed noise.
@@ 4 Numerical Experiments
In this section, we present numerical experiments to evaluate the performance of low-rank Muon (Algorithm 4) and compare our method with vanilla Muon, AdamW, and SGD. The experiments are conducted on GPT-2 pretraining (Section 4.1) and LLaMA pretraining (Section 4.2). All experiments are executed on a server with two NVIDIA A100 GPUs (80 GB). The code to implement the proposed algorithm on these numerical examples is available at https://github.com/dengzhanwang/Low-rank-Muon.
@@ 4.1 GPT-2 Pretraining
In this subsection, we consider pretraining GPT-2 (Radford et al., 2019), a transformer-based language model. We experiment with GPT-2 models of sizes 60M, 135M, 350M, and 1B on the same three datasets tested in the Muon GitHub repository (Jordan et al., ): FineWeb10B, FineWeb100B, and FineWebEdu10B.
@@
We apply low-rank Muon, Muon, SOAP, GaLore, AdamW, and SGD with momentum (SGDM) for GPT-2 pretraining. We implement two low-rank Muon variants with rank parameters 100 and 200, abbreviated as LR-Muon100 and LR-Muon200, respectively. Following the experiments in (Jordan et al., 2024; Lau et al., 2025), we use AdamW to train the embedding and head layers for both Muon and low-rank Muon. We initialize all methods with the same weights from pretrained GPT-2 models, and terminate all methods after one training epoch, consisting of 5000 iterations. We compare all methods using validation perplexity, a standard metric in foundation model training (e.g., (Radford et al., 2019; Touvron et al., 2023; Vaswani et al., 2017)), which is defined as the exponential of the validation loss and serves to amplify performance differences. The algorithmic parameters are carefully tuned to suit each method well in terms of computational performance. More details about the experimental setups, including algorithm and GPT-2 model parameters, are provided in supplementary materials.
@@
We present a comparison of validation perplexity and computational time in Table 1. From this table, we observe that for GPT-2 with a model size of 60M, our low-rank Muon achieves slightly worse validation perplexity than full-rank Muon, but performs better than AdamW, SGDM, SOAP, and GaLore. For GPT-2 models with larger sizes, our low-rank Muon achieves better validation perplexity than all other competing methods. We also observe that the computational time of our low-rank Muon improves upon that of vanilla Muon, although all Muon-type methods remain slower than SGDM and AdamW. These observations show that our low-rank orthogonalization enables obtaining more generalizable GPT-2 models when the model size is large, but may be less effective for smaller models. In addition, our low-rank orthogonalization saves computational time compared to orthogonalization by Newton-Schulz iterations, but the gain is not as significant as those presented in Section 3.1, since the forward and backward passes of neural networks account for the majority of the computational time.
We plot the convergence behavior of validation perplexity versus training iterations in Figure 4. From this figure, we observe that for each model size, the competing methods exhibit similar performance patterns across the three datasets. For the model size of 60M, all Muon-type methods consistently outperform AdamW and SGDM, while our low-rank Muon performs slightly worse than vanilla Muon. For the 135M model, our low-rank Muon outperforms vanilla Muon and other training methods. Vanilla Muon performs similarly to AdamW, and all methods show a large improvement compared to SGDM. For the model size of 350M, our low-rank Muon has slightly better performance than AdamW, while outperforming Muon and SGDM by a larger margin. These observations support that our low-rank Muon improves upon vanilla Muon in GPT-2 training and is more effective for models with larger sizes.